# Beyond the Linear Separability Ceiling

### v1

@@ 5 Linear reasoning bottleneck
Our framework, applied to 8 SoTA VLMs across two datasets and tested with 8 prompt formats each, reveals a widespread linear reasoning bottleneck. As shown in Figure 3, models’ generative performance ($acc_{gen}$) struggles to surpass their LSC. Full results are in Appendix C.
Figure 3: generative accuracy vs. linear separability for all baseline models. The diagonal dashed line is the LSC benchmark ($y=x$). Points above the line indicate that a model’s generative performance has surpassed its LSC. The plot shows that most evaluations struggle to consistently surpass the benchmark, demonstrating a prevalent linear reasoning bottleneck.
Pathways to success and failure. However, some models surpass the LSC, revealing distinct success pathways. For instance, Pixtral 12B and Qwen2.5-VL 72B consistently achieve a generative performance statistically superior to their LSC. In contrast, other models like Gemma3 27B fail to do so reliably. Table 1 highlights the dramatically different mechanisms behind these outcomes.
Table 1: Performance comparison on the OpenWorld dataset, revealing two primary reasoning pathways. The first, used by Pixtral, increases the linear separability of representations ($acc_{final}>LSC$). The second is a post-representation pathway, where Gemma3 27B represents an extreme case operating on embeddings with no linear separability ($acc_{final}\approx 50\%$), while Qwen2.5-VL 72B is an intermediate case with only a partial drop.
Model Direct acc (%) CoT acc (%) LSC ($acc_{vision}$, %) Final rep. ($acc_{final}$, %)
Pixtral 12B $79.4\pm 3.5$ $84.2\pm 3.2$ $76.0\pm 3.7$ $\mathbf{88.0\pm 2.8}$
Gemma3 27B $93.2\pm 2.2$ $87.2\pm 2.9$ $\mathbf{88.6\pm 2.8}$ 50.0
Qwen2.5-VL 72B $93.6\pm 2.1$ $93.2\pm 2.2$ $\mathbf{86.8\pm 3.0}$ $74.8\pm 3.8$
Our analysis reveals two pathways to surpass the LSC. The first, non-linear representation refinement, is exemplified by Pixtral, which transforms its initial embeddings into a more linearly separable format ($acc_{final}>LSC$). The process suggests that next-token prediction itself can function as a representation learner, enhancing separability in a way that shares deep theoretical underpinnings with contrastive learning (Alshammari et al., 2025). While its final generative accuracy is not a direct readout of these embeddings, its success is characterized by a strong statistical alignment between its generative output and the classifications from a linear probe on these improved final representations.
The second observed route is post-representation non-linear reasoning pathway, where the LLM remaps the initial visual embeddings into a more complex representational geometry ($acc_{final}<LSC$). Gemma3 27B is an extreme example of this, where its final embeddings collapse to chance-level separability ($acc_{final}\approx 50\%$) and its performance is brittle across different prompts. In contrast, Qwen2.5-VL 72B provides a stable and highly effective implementation of this pathway. Although its final representation accuracy also drops relative to its LSC, the model achieves consistent, SoTA performance. This suggests Qwen successfully leverages a sophisticated non-linear reasoning mechanism that can operate on these less-separable embeddings. This pathway highlights that a high LSC is not sufficient on its own; success depends on a robust reasoning process that can navigate non-linear representational spaces.
The disconnect between separability and generation. Generally, better linear separability in a model’s vision embeddings (LSC) correlates with better generative performance. A $\chi^{2}$ test shows a statistically significant dependence between the LSC probe and generative performance in 97 out of 128 instances with it being an inverse dependency only for 2 instances. A deeper look, however, reveals a complex disconnect between embedding separability and generative output. While model’s processing attunes its final representations to its generative logic (significant in 103 out of 128 instances), the relationship is often counterintuitive, with 29 instances showing a significant inverse dependence. The model’s reasoning can therefore be disconnected from, or even opposed to, the logic of a linear probe on the same embeddings. While high final separability is not a prerequisite for success, this suggests a hypothesis for creating more robust and interpretable models: train them to improve the linear discriminability of their final embeddings while ensuring that structure is positively aligned with the generative process.
6 Surpassing the LSC
Our baseline analysis in Section 5 revealed a widespread linear reasoning bottleneck where VLMs struggle to surpass their LSC. Through targeted fine-tuning, we systematically probe intervention points within the VLM, from the vision-language interface to the LLM’s reasoning pathways, to pinpoint and resolve the bottleneck.
6.1 Experimental setup
To pinpoint the most effective intervention points, we employ several parameter-efficient fine-tuning (PEFT) and minimal tuning methods. Each is trained with two distinct objectives to assess its impact on generative accuracy and representation quality. Full hyperparameters for all methods are detailed in Appendix D and Appendix F.
    Interface adaption: To test if the LSC can be surpassed by minor adjustments at the modality interface, we tune isolated components: either the vision-language projection layer or an additive vision bias vector on initial patch embeddings.
    Prompt tuning: By adding trainable soft prompts to the input sequence while keeping the VLM’s core weights frozen, we test whether activating dormant reasoning pathways is enough when provided with an optimal input prompt.
    LoRA: We adapt the model by injecting trainable, low-rank matrices into the attention layers of both the vision encoder and language model. This evaluates how adapting core weights can surpass the linear reasoning bottleneck and improve abstract reasoning.
@@ 6.1 Experimental setup
Objectives. We train the PEFT parameters using two objectives. Next-token prediction ($\mathcal{L}_{\text{NT}}$), the standard language modeling objective of maximizing the likelihood of the ground-truth output tokens. And InfoNCE + NTP ($\mathcal{L}_{\text{combined}}$) objective, which augments $\mathcal{L}_{\text{NT}}$ with an explicit contrastive loss ($\mathcal{L}_{\text{sim}}$) by means of a weighted sum.
@@
This contrastive term explicitly encourages the final embeddings of a query image to be closer to its true category centroid than the alternative. The motivation is to directly improve the linear separability of the final representations, thereby raising the separability ceiling for the task while co-optimizing for generative correctness, which is further detailed in Appendix E.
6.2 Pinpointing the location for intervention
To pinpoint the source of performance gains, we conducted a targeted ablation study on the Phi model with the OpenWorld dataset. The results, summarized in Table 2, reveal that the location of the intervention is the critical factor in surpassing the linear reasoning bottleneck.
Table 2: Reported accuracies for the Phi model on the OpenWorld dataset. The ID evaluation uses the interleaved prompt structure seen during training. For the OOD test, we use a labeled prompt structure that groups all images by category at the end of the prompt. In the generative columns, bold indicates the performance is not significantly lower than the LSC, while an underline indicates the LSC is surpassed.
Tuning method
ID (%)
OOD (%)
LSC (%)
Direct baseline $59.0$ $\mathbf{79.4}$ $84.0$
Vision bias ($\mathcal{L}_{\text{NT}}$) $\mathbf{75.0}$ $\mathbf{68.4}$ $73.8$
Vision projector ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{90.2}}$ $73.8$ $84.0$
Postfix ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{94.2}}$ $\underline{\mathbf{90.2}}$ $84.2$
Prompt ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{90.4}}$ $\mathbf{83.4}$ $84.2$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{89.8}$ $\mathbf{88.6}$ $84.2$
LoRA ($\mathcal{L}_{\text{NT}}$ llm-only) $\mathbf{89.8}$ $\mathbf{89.0}$ $84.2$
Structural overfitting.
Minimalist interventions at the vision-language interface, such as vision bias and projector tuning, lead to structural overfitting rather than robust reasoning, as seen in Table 2. They fail to generalize to new prompt formats. Projector tuning’s performance is entirely dependent on prompt syntax, and vision bias tuning succeeds only by degrading the visual representations (lowering the LSC). Since these interface-level adjustments encourage superficial learning, we discarded them and froze the projector for all subsequent experiments.
The bottleneck is reasoning, not perception. Our results pinpoint the bottleneck to the LLM’s reasoning pathways, not the vision encoder’s perception. An ablation study confirmed this (Appendix G), applying LoRA to the vision encoder yielded no performance gain over adapting the LLM alone. This suggests any adaptations within the vision encoder are a form of naïve loss minimization (Prieto et al., 2025), merely scaling existing features rather than learning more discriminative ones. The core issue is therefore not a perceptual deficit but a deeper failure in how the LLM processes its visual inputs.
6.3 Activation vs. Adaptation
With the bottleneck identified in the LLM’s reasoning module, we find that targeted interventions can robustly unlock latent abilities. Our findings (Appendix H) reveal that the required method for unlocking these pathways depends on the reasoning task, highlighting a distinction between two approaches: activating latent skills versus adapting core weights. The success of postfix tuning, a methodological control that cannot alter visual representations because its learnable tokens are placed after the images, provides conclusive evidence that VLMs possess powerful, dormant reasoning pathways. For semantic concepts (OpenWorld), performance was an activation issue; prompt tuning is as effective as LoRA, suggesting that the model’s inherent skills only need to be steered by an optimal input, process similar to prompt engineering (Burns et al., 2023; Brown et al., 2020). For complex relational reasoning (HOI) activation was insufficient. Prompt tuning struggles where LoRA succeeds, demonstrating that these harder tasks require a fundamental adaptation of the model’s internal weights. This reveals a core principle for these models: activating existing skills suffices for simple concepts, but unlocking relational reasoning demands deeper model adaptation.
@@ 6.4 Comparing objectives
Table 3: Summary of PEFT performance on Bongard tasks. Reported accuracies presented as $accuracy\pm ME$. In the generative columns, bold indicates the generative performance is not significantly lower than the LSC, while an underline indicates the LSC is surpassed. Superscripts on separability scores denote a significant dependence (p < 0.05) with the predictions of the corresponding generative method (G).
@@
Phi OpenWorld Direct baseline $59.0\pm 4.3$ $84.0\pm 3.2^{G}$ $76.4\pm 3.7^{-G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{89.8\pm 2.7}$ $84.2\pm 3.2^{G}$ $74.0\pm 3.8$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{96.2\pm 1.7}}$ $84.2\pm 3.2^{G}$ $93.2\pm 2.2^{G}$
HOI Direct baseline $52.1\pm 3.5$ $71.9\pm 3.1$ $60.5\pm 3.4^{-G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{77.4\pm 2.9}}$ $71.1\pm 3.1^{G}$ $61.1\pm 3.4^{G}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{78.9\pm 2.8}}$ $71.8\pm 3.1^{G}$ $79.4\pm 2.8^{G}$
Pixtral OpenWorld Direct baseline $\mathbf{72.4\pm 3.9}$ $76.0\pm 3.7^{G}$ $87.2\pm 2.9^{G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{93.4\pm 2.2}}$ $76.6\pm 3.7$ $87.2\pm 2.9^{G}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{95.0\pm 1.9}}$ $74.4\pm 3.8$ $96.2\pm 1.7^{G}$
HOI Direct baseline $\mathbf{57.8\pm 3.4}$ $62.7\pm 3.4^{G}$ $70.2\pm 3.2^{G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{78.0\pm 2.9}}$ $61.6\pm 3.4^{G}$ $74.9\pm 3.0^{G}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{79.6\pm 2.8}}$ $63.1\pm 3.3^{G}$ $77.8\pm 2.9^{G}$
Gemma3 4B OpenWorld Direct baseline $76.0\pm 3.7$ $89.8\pm 2.7$ $50.0\pm 4.4^{G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{92.4\pm 2.3}$ $89.8\pm 2.7^{G}$ $50.0\pm 4.4$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{95.6\pm 1.8}}$ $89.8\pm 2.7$ $96.6\pm 1.6^{G}$
HOI Direct baseline $56.5\pm 3.4$ $74.1\pm 3.0$ $50.0\pm 4.4^{G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{84.2\pm 2.5}}$ $74.1\pm 3.0^{G}$ $50.0\pm 4.4$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{84.2\pm 2.5}}$ $74.1\pm 3.0^{G}$ $83.2\pm 2.6^{G}$
Generative performance surpassing separability ceiling. Fine-tuning with the $\mathcal{L}_{\text{NT}}$ objective allows generative accuracy to significantly surpass the LSC (Table 3). This occurs because the generative model can implement a complex, non-linear decision boundary that our linear separability probe cannot. The LSC is therefore not an absolute performance limit but a benchmark for linear reasoning, which models can learn to bypass by activating more sophisticated classification pathways.
Modality alignment and raising the separability ceiling. While the $\mathcal{L}_{\text{NT}}$ objective fosters post-representation non-linear reasoning pathways that bypass the LSC, the combined objective ($\mathcal{L}_{\text{combined}}$) works by directly improving the embedding space itself through non-linear representation refinement. Its contrastive loss term ($\mathcal{L}_{\text{sim}}$) explicitly optimizes the final representations to be more linearly separable, which raises the representation accuracy and aligns the generative process with this improved structure. This alignment is confirmed by the consistent, significant dependence between generative outcomes and final-layer similarity predictions for all LoRA models trained with $\mathcal{L}_{\text{combined}}$ (Table 3), a consistency the $\mathcal{L}_{\text{NT}}$ objective fails to achieve.
This convergence of generative and linear probe accuracy indicates the performance bottleneck shifts from reasoning back to perception. The primary barrier is now the quality of the visual representations themselves. Given that applying LoRA to adapt the vision encoder’s feature extraction process yielded no improvement to the LSC, pointing to a more fundamental limitation. This suggests the remaining perception issue is likely architectural and cannot be resolved by further training with the methods and objectives explored in this study.
Both objectives offer largely equivalent performance on relational HOI task, with $\mathcal{L}_{\text{combined}}$ being better on semantic OpenWorld task. For a visual analysis of how each objective alters the models’ internal attention mechanisms to produce these outcomes, see Appendix M.
6.5 Measuring generalization
@@
Conceptual Generalization. The performance gains from fine-tuning stem from genuine conceptual understanding rather than memorization. This is demonstrated on the Bongard-HOI dataset, where our models show robust generalization by maintaining high accuracy on test splits with entirely unseen objects and actions (Appendix I). In fact, top models often achieve their peak performance on these most challenging splits. This finding is reinforced by strong results on the OpenWorld test set, which consists entirely of new concepts by design.
Structural generalization. Beyond conceptual understanding, we evaluated for structural generalization by testing model robustness to changes in prompt format (Appendix H). We found a critical trade-off: while the standard $\mathcal{L}_{\text{NT}}$ objective yields structurally robust models, the $\mathcal{L}_{\text{combined}}$ objective, which improves representation quality, can be brittle. This fragility is evident on the HOI dataset, where models trained with the combined objective fail on out-of-distribution (OOD) prompts despite producing highly separable final-layer representations. This indicates the generative pathway has overfit to the prompt’s syntax, failing to interpret its own well-structured embeddings in a new format. The superior OOD performance of postfix tuning over methods that include a prefix further confirms this sensitivity to prompt structure.
Generalization under domain shift. A rigorous cross-domain evaluation revealed that successful generalization depends on aligning the fine-tuning objective with the model’s intrinsic reasoning strategy (Appendix J). For models like Pixtral and Phi, which naturally excel at representation refinement, the $\mathcal{L}_{\text{combined}}$ objective, which explicitly enhances representations, proved most resilient. Remarkably, when trained on OpenWorld, Pixtral’s generative performance on HOI (71.0%) also surpassed its LSC (63.1%), showing that its learned skill had successfully transferred. Conversely, for Gemma3, which relies on a post-representation pathway, the standard $\mathcal{L}_{\text{NT}}$ objective provided more robust (though still limited) generalization. This highlights that true generalization is not achieved by a single best method, but by matching the intervention to the model’s architectural strengths.

### v3

@@ 5 Finding: A pervasive alignment gap
Our framework, applied to 8 SoTA VLMs across two datasets and tested with 8 prompt formats each, reveals a pervasive alignment gap. As shown in Figure 3, a model’s generative performance is highly variable and frequently fails to surpass its own LSC. We identify this widespread issue as an alignment gap, where a model’s reasoning pathways are not effectively aligned with its own visual representations. Full results are in Appendix F.
Figure 3: Vertical axis is the generative performance across different prompts, horizontal axis is the linear probe classification accuracy. The datapoints in the green region are the instances when the model is successfully using its reasoning pathways to generatively classify the query image significantly better than a linear probe on inputs would, surpassing the linear separability ceiling. In the red region, however, alignment gap persists and the non-linear nature of VLM is failing to outperform a linear classifier. Statistical comparison accounts for 95% confidence intervals of both metrics.
5.1 Analysis of LSC-surpassing reasoning pathways
While most configurations are constrained by this bottleneck, a few successful models reveal how it can be overcome. Our analysis shows that successful models channel the transformer’s iterative refinement process into two distinct computational strategies. These reasoning pathways are distinguished by the geometric transformations applied to the visual embeddings, with Table 1 highlighting the divergent outcomes.
Table 1: Performance metrics for selected SoTA models on the OpenWorld dataset, comparing direct and CoT generative accuracy with initial (LSC) and final representation linear separability accuracies.
Model Direct acc (%) CoT acc (%) LSC (%) Linear separability (final, %)
Pixtral 12B $79.4$ $84.2$ $76.0$ $\mathbf{88.0}$
Gemma3 27B $93.2$ $87.2$ $\mathbf{88.6}$ collapsed (50.0)
Qwen2.5-VL 72B $93.6$ $93.2$ $\mathbf{86.8}$ $74.8$
Human $91.0$ - - -
Enhancing linear separability. The first and rarer strategy, employed only by Pixtral, is to enhance linear separability using non-linear processes. This effectively continues the bottom-up perceptual work of the vision encoder. Its reasoning pathway processes the initial embeddings into a more linearly separable format. It refines features in a manner that aligns with the principles of contrastive learning (alshammari2025icon). This results in a strong statistical alignment between its generative output and a linear probe on the improved final representations.
Non-linear decision logic. The vast majority of successes, however, were achieved by implementing a non-linear decision logic, a form of top-down reasoning that computes a solution not linearly readable from the visual embeddings themselves. It leverages the computational depth of the model to form a non-linear decision boundary. Gemma3 represents an extreme case, where its final embeddings undergo a representational collapse with respect to linear separability (accuracy is 50%). This pathway highlights that a high LSC is a valuable starting point, but success also depends on a powerful non-linear process operating within a non-linear representational geometry.
Table 2: Summary of $\chi^{2}$ statistical dependence tests between generative performance and linear probe outcomes across 128 VLM configurations, where (+) and (-) denote significant positive/inverse dependence (both methods tend to agree/disagree), and (Ø) for no significant dependence.
Probe location + - Ø
Vision embeddings 95 2 31
Final embeddings 74 29 25
Statistical dependence of these pathways. To better understand the connection between these reasoning pathways, we perform a $\chi^{2}$ test for statistical dependence between the generative predictions and the outcomes of our linear probes (Table 2). The generative output is positively correlated with linear separability probe on most models’ vision and final embeddings, suggesting that more linearly separable representations generally lead to better generative performance. Interestingly, generative performance is inversely correlated with the linear separability of the final embeddings, implying the model is making a guess different from what it perceives to be the right guess. To move beyond this correlational ambiguity and identify the true drivers of performance, we first investigate the optimal intervention points for enhancing the model’s generative reasoning, and then further explore this behavior.
5.2 Enhancing generative accuracy beyond LSC
To surpass the LSC, we compare three intervention types: at the vision-language projector, input-level guidance, and full model adaptation. Input guidance includes prompt tuning (trainable soft prompts with a meta-learning objective). Full adaptation uses LoRA, injecting trainable matrices into attention and MLP layers. We also test robustness to unseen prompt formats, with full details in Appendix G.
Table 3: Phi model accuracies (%) on the OpenWorld dataset. The in-distribution (ID) evaluation uses the interleaved prompt structure seen during training. For the out-of-distribution (OOD) test, we use a labeled prompt structure that groups all images by category at the end of the prompt. In the generative columns, bold indicates the LSC is surpassed.
Tuning method
ID
OOD
LSC
Direct baseline $59.0$ $79.4$ $84.0$
Projector ($\mathcal{L}_{\text{NT}}$) $\mathbf{90.2}$ $73.8$ $84.0$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $\mathbf{94.2}$ $\mathbf{90.2}$ $84.2$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $\mathbf{90.4}$ $83.4$ $84.2$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{92.2}$ $\mathbf{89.8}$ $84.4$
LoRA ($\mathcal{L}_{\text{NT}}$ llm-only) $\mathbf{92.4}$ $\mathbf{89.6}$ $84.2$
Pinpointing the location for intervention. An ablation study revealed that the intervention’s location is key for surpassing the alignment gap. The results, summarized in Table 3, show that minimalist interventions at the vision-language interface lead to prompt overfitting rather than robust reasoning. We therefore froze the projector in all subsequent experiments. Our results pinpoint the bottleneck is in the LLM’s reasoning pathways, not the vision encoder’s perception. An ablation study confirmed this (Appendix J), as applying LoRA to the vision encoder resulted in nearly identical predictions and no additional performance gain compared to adapting the LLM alone. This suggests any adaptations within the vision encoder are likely a form of naïve loss minimization (prieto2025grokking), merely scaling existing features rather than learning more discriminative ones. The core issue is therefore not a perceptual deficit but a deeper failure in how the LLM processes its visual inputs.
Activation vs. adaptation. With the bottleneck identified in the LLM’s reasoning module, we find that targeted interventions can robustly unlock latent abilities. Our findings reveal a critical distinction between two approaches: activating latent skills versus adapting core weights, with the required method depending on the reasoning task. The success of postfix tuning, a methodological control that cannot alter visual representations, provides evidence that VLMs possess powerful, dormant reasoning pathways capable of performing non-linear decision logic, activating reasoning pathways.
Table 4: Reported accuracies (%) for Phi model on both datasets, where bold indicates the LSC is surpassed.
Method OpenWorld HOI
Direct baseline $59.0$ $52.1$
LSC $84.0$ $71.9$
Postfix tuning $\mathbf{94.2}$ $63.2$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{92.2}$ $\mathbf{78.6}$
As shown in Table 4, performance on OpenWorld is an activation issue. Prompt-based methods are effective, achieving performance comparable or even superior to LoRA. This suggests the model’s inherent skills only need to be steered by an optimal input, a process similar to prompt engineering (burns2023discovering; brown2020language). On the relational HOI task, activation methods prove insufficient, requiring deeper adaptation via LoRA to substantially boost performance. This reveals a core principle, holding true for all tested models (Appendix K), that activation suffices for atomic semantic concept comparison while adaptation is required for reasoning over relational semantic concepts.
6 Connecting the two reasoning pathways
Our findings so far show that interventions using the standard $\mathcal{L}_{\text{NT}}$ objective improve the non-linear decision logic pathway to surpass the LSC. This raises a question, however, what happens if we combine both reasoning pathways? Instead of forcing a non-linear decision logic to overcome poor separability, can we redirect these non-linear computational pathways to improve the linear separability of final embeddings also?
@@ 6 Connecting the two reasoning pathways
To test this, we introduce a combined training objective designed to do exactly that, consisting of two objectives: next-token prediction ($\mathcal{L}_{\text{NT}}$), the standard language modeling objective of maximizing the likelihood of the ground-truth output tokens; and an explicit contrastive loss ($\mathcal{L}_{\text{sim}}$) designed to identify a positive sample from a set of negative samples (oord2018representation) by means of a weighted sum.
@@
This contrastive term explicitly encourages the final embeddings of a query image to be closer to its true category centroid than the alternative, as seen on Figure 4, detailed further in Appendix H.
(a) Embedding space before.
(b) Embedding space after.
Figure 4: Principal component analysis (PCA) visualization of the embedding space before (a) and after (b) applying the contrastive term. As PCA applies only linear transformations, it preserves the relative geometric structure of the original space. In the plots, the green cross ($\mathbf{+}$) is the query image’s mean embedding, while the larger blue and red ($\mathbf{X}$) markers are the centroids for the positive and negative image sets, respectively. Before training, representations for positive (blue dots), negative (red dots), and query (green dots) images are heavily overlapped. After training with the contrastive loss, the space is restructured: the query image’s representations (green dots) now occupy the same geometric region as the positive set (blue dots), clearly separating from the negative set (red dots).
We find that the balance between these objectives is critical; as shown in our sensitivity analysis in Appendix N, a dynamic dual cosine schedule results in a more controlled dynamic between the two loss counterparts.
@@ 6.1 Comparing objectives
Table 5: Summary of PEFT performance on Bongard tasks. In the generative columns, bold indicates the LSC is surpassed. Superscripts on separability scores denote a significant dependence (p < 0.05) with the predictions of the corresponding generative method (G).
@@
Phi OpenWorld Direct baseline $59.0$ $84.0^{G}$ $76.4^{-G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{92.2}$ $84.4^{G}$ $74.8$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{95.6}$ $84.4^{G}$ $93.8^{G}$
HOI Direct baseline $52.1$ $71.9$ $60.5^{-G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{78.6}$ $71.9^{G}$ $63.6^{G}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{79.2}$ $71.9^{G}$ $82.0^{G}$
Pixtral OpenWorld Direct baseline $72.4$ $76.0^{G}$ $87.2^{G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{93.4}$ $76.6$ $87.2^{G}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{95.0}$ $74.4$ $96.2^{G}$
HOI Direct baseline $57.8$ $62.7^{G}$ $70.2^{G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{78.0}$ $61.6^{G}$ $74.9^{G}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{79.6}$ $63.1^{G}$ $77.8^{G}$
Gemma3 4B OpenWorld Direct baseline $76.0$ $89.8$ $\text{collapsed (50.0)}^{G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $92.4$ $89.8^{G}$ collapsed (50.0)
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{95.6}$ $89.8$ $96.6^{G}$
HOI Direct baseline $56.5$ $74.1$ $\text{collapsed (50.0)}^{G}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{84.2}$ $74.1^{G}$ collapsed (50.0)
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{84.2}$ $74.1^{G}$ $83.2^{G}$
Human OpenWorld 91.0 - -
HOI 91.4 - -
Aligned representational geometry. Training with $\mathcal{L}_{\text{combined}}$ induces a significant and consistent statistical dependence between the generative and the final-layer linear probe predictions. Furthermore, their respective accuracies also converge. This convergence indicates a shift in the model’s computational strategy. Its non-linear pathways are re-directed: instead of executing a non-linear decision logic on top of the representations, they are repurposed to actively refine the representations into a more linearly separable format based on surrounding context. We observe evidence for this in the model’s internal attention mechanisms, where the most significant differences are concentrated in the later layers. This finding is consistent with research suggesting that reasoning mainly happens in middle-to-late layers, whereas perception is encoded in early layers (chen2025bring). A visual analysis of these attention changes is available in Appendix P.
@@
Conceptual Generalization. The performance gains from fine-tuning stem from genuine conceptual understanding rather than memorization. This is demonstrated on the Bongard-HOI dataset, where our models show robust generalization by maintaining high accuracy on test splits with entirely unseen objects and actions (Appendix L). This finding is reinforced by strong results on the OpenWorld test set, which consists entirely of new concepts by design. Whether further HOI progress needs better visual features or smarter reasoning pathways remains unclear.
Model Method Acc. (%)
CLIP (ViT-L/14) Baseline 27.75
Phi Baseline 16.75
$\mathcal{L}_{\text{NT}}$ 18.25
$\mathcal{L}_{\text{combined}}$ 29.25
Pixtral Baseline 28.75
$\mathcal{L}_{\text{NT}}$ 43.50
$\mathcal{L}_{\text{combined}}$ 54.75
Gemma3 4B Baseline 5.25
$\mathcal{L}_{\text{NT}}$ 5.75
$\mathcal{L}_{\text{combined}}$ 12.25
Table 6: Text retrieval on Winoground, comparing baselines with HOI-trained LoRAs. The CLIP model is included as an LSC baseline for the Phi model.
Generalization under domain shift. Cross-domain evaluations reveal that the nature of the skill learned during fine-tuning is a key determinant of its transferability. For instance, the relational reasoning skills acquired from the HOI dataset transferred broadly across both our cross-Bongard task and the Winoground benchmark, whereas the atomic concept comparison skills from the OpenWorld dataset proved less generalizable (Appendix M). This distinction is particularly evident on Winoground, where only the HOI-trained models achieved substantial performance gains (Table 6). Here, our $\mathcal{L}_{\text{combined}}$ objective achieved the most successful transfer by consistently improving text retrieval accuracy for all models, a success we attribute to the inherent need for inter-image comparison in Bongard-style problems. Furthermore, this out-of-domain evaluation validated our LSC framework by demonstrating both its broader applicability and the difficulty of surpassing the LSC, which a fine-tuned Phi model could not significantly surpass.
7 Discussion, limitations and future-work
Discussion. Recent theoretical work establishes that transformer language models are almost surely injective, meaning they are structurally lossless and different inputs provably map to different internal representations (injective2025). The complete input information is therefore preserved. However, this injectivity is highly non-linear; the input undergoes a deep composition of attention, normalization, and activation functions. This creates a gap: while all information is likely present, it is not necessarily accessible. From a computational perspective, a linear representation is more likely to generalize to downstream tasks as it is more amenable to information extraction.
Figure 5: Manifold structure.
However, models are typically trained with a next-token prediction loss only, which does not explicitly enforce linear separability. Consequently, when models are subjected to instruction-tuning or reinforcement learning, the task-specific learning signal can cause representational degradation, where linearly decodable information is diminished. The LSC framework detects this degradation, while the auxiliary contrastive signal actively restores and promotes linear separability. As visualized in Figure 5, this signal restructures the manifold into one-dimensionally linear “rays” (further detailed in Appendix B), effectively structuring representations into globally consistent semantic directions.
Limitations. While $\mathcal{L}_{\text{combined}}$ objective successfully restructures vision representations, it appears to compromise the complex, non-linear decision logic built upon them, effectively reintroducing an image-to-text alignment gap. This brittleness became apparent when testing model robustness on the HOI dataset (Appendix K). Notably, all three models trained with $\mathcal{L}_{\text{combined}}$ suffered performance degradation when the prompt format was altered, whereas only one model trained with $\mathcal{L}_{\text{NT}}$ objective exhibited a similar decline. This pattern extends to multiple-choice VQA benchmarks, where models trained with the $\mathcal{L}_{\text{combined}}$ objective experienced steeper performance drops compared to their $\mathcal{L}_{\text{NT}}$ counterparts (Appendix A).
Future work. It is inherent that significantly altering the geometry of representations disrupts the reasoning pathways calibrated to them. To mitigate this, we propose a promising avenue for future work: enforcing image-to-text alignment where the one-dimensional linearity observed in our visual manifold is also explicitly promoted within text representations. By anchoring semantically analogous visual embeddings to this structured textual space, we can resolve the “split-brain” pathology of current VLMs — characterized by disjoint heuristics for visual and textual processing — thereby establishing a foundation for unified, truly multimodal reasoning.